### v1

# Single-beam grating-chip 3D and 1D optical lattices

Abstract
Ultracold atoms are crucial for unlocking truly precise and accurate quantum metrology, and provide an essential platform for quantum computing, communication and memories. One of the largest ongoing challenges is the miniaturization of these quantum devices. Here, we show that the typically macroscopic optical lattice architecture at the heart of many ultra-precise quantum technologies can be realized with a single input laser beam on the same diffractive chip already used to create the ultracold atoms. Moreover, this inherently ultra-stable platform enables access to a plethora of new lattice dimensionalities and geometries, ideally suited for the design of high-accuracy, portable quantum devices.
Optical lattices, formed by the standing wave of two or more interfering laser beams are an essential platform for many quantum optics experiments [1, 2, 3]. The extremely long atomic coherence times possible in such optical traps have been used extensively in some of the most sensitive measurements to date, notably extending the lifetime in atomic memories many orders of magnitude to several seconds [4], and allowing optical lattice clocks [5] to reach fractional frequency accuracies at the level [6, 7]. Such clocks are typically built on 1D lattices, but 3D [8, 9] and 2D lattices [10] can also be used.
Lattices in one [11], two [12, 13], and three [14] dimensions were also vital components for many key experiments with quantum degenerate gases. By selectively preparing a single 2D layer from such 3D lattices, quantum gas microscopes can be prepared in bosonic [15, 16] and fermionic [17, 18, 19] atoms, in an increasing range of atomic species [20, 21, 22, 23]. Optical lattices for storage, augmented by optical tweezers for transport [24], can also be used in the burgeoning field of neutral-atom quantum computing [25, 26].
In recent years there has been a concerted effort to develop portable cold-atom sensors, translating the excellent performance of lab-based experiments to real-world applications. Optical-lattice-based devices utilise laser-cooled atomic clouds, created in magneto-optical traps (MOTs [27]) using overlapping near-resonant laser beams, which are subsequently loaded into the lattice potentials. The lattices themselves are typically generated from a single laser at a far-detuned wavelength which is then split into multiple intersecting laser paths 111Often with frequency shifts between non-colinear beams., adding more complexity to beam delivery. Mirrors aid stability in 1D lattices, by providing a fixed node of the optical standing wave, however this stable point is often many centimetres away from the atoms. Multiple intersecting 1D lattices can generate higher-dimensional lattices, if they are carefully mutually orthogonally polarised and/or frequency shifted to avoid optical interference which can shake the lattice or even change its geometry [29].
Previous work on atom chips has demonstrated atomic loading from a MOT into a 1D optical lattice by way of a magnetic surface trap [30, 31], however the mirror and 3D MOTs used require four and six beams, respectively. The integration of a single-beam pyramid MOT [32] and 1D and 2D optical lattices [33] has also been demonstrated, but the lattice did not use the same input beam direction as the MOT, and lattice mirrors were mm from the atoms. The equivalent to a single-input-beam MOT with a 1D lattice along the same input axis has been achieved in interferometry [34], however the lattice in this case is used as a beamsplitter instead of storage, and atom number may be limited [35].
Grating magneto-optical traps (GMOTs) offer a convenient platform for the formation of ultracold atomic clouds with a single input trapping beam [36, 37]. This technology has now been extended to a range of atomic species [38, 39, 40, 41] and applications, and is well-placed for adoption in compact, cold-atom quantum sensors [42, 43, 44]. Here we extend this single-input beam architecture to the formation of optical lattices [36]. By aligning a single high intensity far-detuned laser beam onto different areas of the grating chip, 1D, 2D and 3D lattices can be formed – on the same grating used for the GMOT. All lattice beams share a common nodal plane, pinned to the grating chip surface mm from the atoms, so the resulting optical lattice is stable against the drift often seen in macroscopic systems. We anticipate that this simple, highly compact and stable optical lattice geometry will be ideal for many portable quantum-optic experiments.
Fig. 1 a illustrates the central of the GMOT chip we use to create a 3D lattice. The chip comprises three sectors of one-dimensional diffraction grating (gray triangles). The binary grating sectors have properties and fabrication methods as detailed in [36, 45, 46, 41], and are optimised for the atomic species utilised here – 87Rb. We use a similar theoretical description to our MOT modelling papers [35, 47], however here we consider the coherent superposition of electric fields rather than the Doppler- and Zeeman-effect sensitive sums of intensity-based average radiation pressures (see Supplementary Material (SM)). Theoretical cross-sections of the 3D lattice itself are shown in Fig. 1 a and b, illustrating the different lattice regions with their distinct geometries and dimensions. The 1D lattice model is shown in Fig. 1 c, with experimental data in d.
To make the 3D (1D) optical lattices in this paper, we use a single circularly- (linearly-) polarised laser beam, normally incident near the centre of the grating with wavelength nm (nm), power W (W), and intensity radius mm (m). For the near-resonant lattice at nm, the laser red-detuning was GHz relative to the center of the four Doppler-broadened D2 transitions. Our Al-coated grating, has a material reflectivity of 88% at nm and 95% at nm [48]. Due to the grating period nm, the nm light diffracts (-order) at an angle of to the grating normal with diffraction efficiencies into the first and zeroth orders of and respectively [46]. The nm light has most power in the zeroth order , and the first-order diffraction angle is large (82∘) so these beams have no impact on the 1D lattice.
As a basis for loading the optical lattices, 8 87Rb atoms are loaded into a GMOT and cooled to 3K using a standard red molasses cooling stage [36, 37]. Stray DC magnetic fields are cancelled by three pairs of orthogonal Helmholtz coils. The MOT location with respect to the chip, and its overlap with the optical lattice, can be controlled within the GMOT’s optical overlap volume by varying the zero position of the quadrupole magnetic field. The typical MOT location and size are also shown in Fig. 1 a and b.
The grating chip is bonded to the open end facet of a rectangular cuvette vacuum cell using UHV compatible epoxy, allowing a base vacuum pressure at the mbar level, as measured by the ion pump. No bowing of the chip is observed due to pressure differential [49]. The decision to mount the chip in-vacuum was made to allow the lattice to form as close to the chip surface as possible with no glass interface, enabling future use of atoms in chip-surface traps 222The lattices here could in principle be achieved with an ex-vacuo grating chip, near a thin-walled glass chamber, but this could reduce the lattice capture efficiency..
Atom detection is achieved via fluorescence imaging onto a CCD camera sensor. The imaging system is aligned orthogonal to the directions of both the MOT/lattice beam and gravity. A spatially selective focal plane is utilised to reduce background noise originating from scattered light 333Additionally, a notch filter with a peak transmission centred around nm is used when investigating the 1D nm lattice. For both the 3D and 1D lattice geometries s long resonant fluorescent images are taken with the MOT light after a ms time of flight following the extinction of the lattice beam. A calculation of atomic motion during imaging (see SM) quantitatively matches the large measured width of the cloud seen in Fig. 1 d. In both the 3D and 1D lattices, the lattice beam power can be rapidly adjusted by a single-pass acousto-optic modulator (AOM) designed for the beam wavelength.
The nm 3D lattice beam is from a near-resonant diode laser [52] with a tapered amplifier. The broadband amplified spontaneous emission (ASE) pedestal associated with this laser produces some near-resonant light which can heat atoms from the lattice via scattering. To reduce this effect we filter the lattice beam through a cm long C Rb vapour cell [53]. The lattice beam is combined with the MOT beam on a polarising beam splitter. Atoms are loaded into the 3D lattice potentials by switching the lattice beam on at the end of the red molasses sequence, to mitigate the effect of light shifts on the cooling process.
To form the far-detuned 1D lattice a nm ytterbium laser is used. The vertically polarised light is combined with the MOT light on a dichroic mirror and aligned onto the grating chip. Atoms are loaded into the 1D lattice by switching the lattice beam on at the beginning of the red molasses cooling stage.
A log plot of the number of trapped atoms as a function of hold time in both the 3D and 1D lattice is shown in Fig. 1 e. Exponential decays are fitted, excluding data where ms, to eliminate untrapped atoms ejected from the imaging region over this time scale due to gravity. The 1D lattice has a much longer lifetime, this is because it is much deeper and the far-off resonance light also induces less scattering and thereby heating than the near-resonant 3D lattice. The overlap volume between the 3D lattice and the MOT is much larger than for the 1D lattice, which can be seen by the difference in initial lattice atom number.
From the time constant of the fits we estimate a 3D lattice lifetime of ms. Due to the limited power available from the laser a relatively low detuning from the Rb D2 line of GHz was used. As a consequence, atomic heating due to photon scattering is significant, estimated to be kHz for each photon absorption-emission cycle within the deepest (and brightest) parts of the lattice, where the expected potential depth at lattice sites is K. This leads to a maximum heating rate at lattice sites of K/s in good agreement with the observed ensemble heating rate across the lattice of K/s 444Alan Bregazzi, PhD thesis (Strathclyde, 2024).. Along with any residual unfiltered laser ASE, and the relatively shallow lattice depth, this heating accounts for the shorter near-resonant lattice lifetime.
The far-off resonance 1D lattice is expected to have less photon scattering (Hz) with a greater depth at the lattice sites (K). Fitting an exponential decay to lattice atom number data indicates a longer trap lifetime of ms. An estimation of the background vacuum pressure can be made using the known relation between MOT lifetime and vacuum pressure [55, 56, 57]. The shallower potential of the 1D lattice compared to the MOT leads to a shorter background-limited lifetime [37, 58]. When the Rb partial pressure was changed, the MOT fill time was consistently twice as long as the 1D lattice decay time, indicating that both the MOT and 1D lattice are background-pressure limited.
The long hold time of the atoms, far in excess of that possible if they were in free-fall, is strong evidence of the formation of an optical dipole trap. The observed behaviour could not arise due to optical molasses or any light scattering forces. Even if we overestimate the effects, by considering a single laser beam scattering at kHz against gravity, the maximum acceleration is m/s2. Similarly, using a 1D Doppler model with co-propagating beams in the gravity direction with our laser intensity and detuning, the overestimated molasses acceleration reached is a millionth of that due to gravity.
Although direct imaging of the lattice is not possible without a suitable microscope for site-resolved imaging, additional evidence can be obtained via investigating parametric resonances at the lattice trapping frequencies [59, 58]. This resonance technique relies on modulating the lattice potential. When the modulation frequency is equal to twice the trapping frequency , the kinetic energy of the atoms increases exponentially, resulting in parametric excitation to higher vibrational states of the trapping potential. Eventually atoms are ejected entirely from the trapping potential. This process can be easily observed experimentally by an increase in cloud size, due to atoms entering higher vibrational states, or a drop in atom number as atoms are ejected from the trap. In addition to the fundamental parametric resonance at , sub-harmonics can also be excited at frequencies given by , where is an integer [59].
We modulated the lattice beam power in the range using the AOM. In the 3D lattice, modulation is applied for ms after first holding the atoms in the static lattice for ms to allow for free evaporation [1]. The atoms are then imaged after a ms time of flight. In the 1D lattice case, modulation is applied for ms due to the improved lattice lifetime and the need to eject a significant number of atoms from the deeper 1D potentials in order to observe the trap resonances. In all cases the modulation scans are randomised in frequency to negate any effects due to long-term drifts in experimental conditions during the data campaigns.
A spectrum of the cloud spatial standard deviation after modulation, as a function of modulation frequency, illustrates the heating of the atoms in the near-resonant 3D lattice (Fig. 2 a). We see a clear primary peak at around kHz, with a partially-resolved secondary feature at around kHz. Because the 3D lattice is relatively shallow, broad resonances are expected as the atoms explore the full sinusoidal potential, not just the harmonic lattice minima.
We have also determined the expected theoretical trap frequencies in the 3D lattice, allowing for the spread due to lattice intensity variation across the atomic cloud from the Gaussian nature of the input and diffracted beams. For the conditions in Fig. 2 a this yields trap frequencies in the range kHz and kHz, where standard deviations are given, also allowing for the spatially varying lattice population due to the initial MOT cloud distribution used for loading. In addition, the atoms after optical molasses are usually populated uniformly across the magnetic hyperfine levels. While subject to the circularly polarised lattice beam 555Required for balanced diffraction from all three sectors of the grating. the atoms are subject to the theoretical trap frequencies above corrected by a factor depending on which of the five different states they are in.
The experimental 3D lattice heating frequencies therefore match surprisingly well with the expected trap frequencies 666Similar discrepancies between theory and experiment have been reported previously, with experimental frequencies (60-90)% below the maximum theoretical expectations in several separate experiments Ref. [58].. We also found that changing the lattice depth affects the resolution of the primary and secondary features in Fig. 2 a, with data for a range of five powers in the SM. This may be due to atom loss from lower depth lattice sites at lower laser power leading to a remaining trapped atom distribution predominantly occupying the central portion of the lattice structure with a lower spread of trap frequencies.
To further test that the atoms experience a lattice arising due to the dipole force, we measured the primary lattice resonance frequency for a variety of beam powers , and red detunings . By plotting the primary observed resonance frequency as a function of , resonance data for different 3D lattice trap depths can be summarised in a single plot, as shown in Fig. 2b. As expected, these data follow a relation, where here the fitted value is kHzGHz/mW, in good agreement with theory.
The measured parametric resonance frequencies in the 1D lattice (Fig. 2 c) are kHz and kHz, which, like in the 3D case, match well with sub-harmonics of the expected trap frequencies, kHz and kHz. There should be much less spatial variation in trap frequency due to lattice structure in the 1D lattice compared to the 3D lattice, and the atoms are expected to be in the deeper harmonic parts of the lattice, with no trap frequency variation with state. In the 1D case, however, the signal-to-noise of the parametric heating is low, and the (relatively broad) resonances can only be observed via atom number loss. This is due to lower spatial overlap of the 1D lattice with the MOT and hence lower initial atom number (Fig. 1 e), as well as rapid radial expansion during fluorescence imaging (Fig. 1 c, d and the SM). The 1D lattice beam focal position on the grating surface was optimised for atom number, and given the lattice beam Rayleigh range of mm, the beam waist was constant over the entire lattice.
In this paper we have demonstrated optical lattices in one and three dimensions. However, we note that by loading the MOT atoms into a magnetic and/or optical trap, they can be transported beyond the MOT capture volume, and thus future single-input-beam two-dimensional lattices are within easy reach (Fig. 1 a, b).
We have demonstrated experimentally that the GMOT architecture can be extended from a single-input-beam MOT, to both three- and one-dimensional single-input-beam optical lattices on the same chip that is also likely to enable 2D lattices. Because the grating surface, from which all other lattice beams originate, is a node of the electric field the associated beam phases and thereby grating lattices are expected to be much more stable than standard optical lattices derived from multiple centimetre-distant mirrors.
The 3D and 2D lattices arising from our binary grating chip with three-fold symmetry (‘tri’ chips) naturally generate lattices of a topical [62, 63, 64] triangular nature. However, equivalent 3D body-centred-cubic lattices (comprising offset 2D square lattice layers) can be generated using binary gratings based on either four sectors of linear grating (‘quad’ chips [36, 65, 57]), or the larger capture volume 2D checkerboard grating option [36]. By moving beyond binary gratings into multi-level gratings, it is expected that more exotic lattice geometries may also be explored [66, 67, 68]. Given that typical optical lattices have beam waists around m, it seems natural to have a GMOT grating region bordered by a wide variety of small grating patterns that the MOT can be transported to, to easily explore different lattice geometries within the same experimental setup.
We therefore expect our technology is ideally suited for application in a wide range of stable, high-accuracy, portable quantum devices.
Acknowledgements.
The authors would like to thank Arthur La Rooij and Oliver Burrow for valuable discussions. A.B. was supported by a Ph.D. studentship from the Defence Science and Technology Laboratory (Dstl). J.P.M. gratefully acknowledges support from a Royal Academy of Engineering Research Fellowship. We thank the UK EPSRC for funding via grant EP/T001046/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising from this submission. The dataset is available here 777TBD.References
- Grimm et al. [2000] R. Grimm, M. Weidemüller, and Y. B. Ovchinnikov, Optical dipole traps for neutral atoms, in Advances In Atomic, Molecular, and Optical Physics (Elsevier, 2000) pp. 95–170.
- Bloch et al. [2008] I. Bloch, J. Dalibard, and W. Zwerger, Many-body physics with ultracold gases, Reviews of Modern Physics 80, 885 (2008).
- Ludlow et al. [2015] A. D. Ludlow, M. M. Boyd, J. Ye, E. Peik, and P.-O. Schmidt, Optical atomic clocks, Reviews of Modern Physics 87, 637 (2015).
- Dudin et al. [2013] Y. O. Dudin, L. Li, and A. Kuzmich, Light storage on the time scale of a minute, Physical Review A 87, 031801(R) (2013).
- Takamoto et al. [2005] M. Takamoto, F.-L. Hong, R. Higashi, and H. Katori, An optical lattice clock, Nature 435, 321 (2005).
- Takamoto et al. [2020] M. Takamoto, I. Ushijima, N. Ohmae, T. Yahagi, K. Kokado, H. Shinkai, and H. Katori, Test of general relativity by a pair of transportable optical lattice clocks, Nature Photonics 14, 411 (2020).
- Bothwell et al. [2022] T. Bothwell, C. J. Kennedy, A. Aeppli, D. Kedar, J. M. Robinson, E. Oelker, A. Staron, and J. Ye, Resolving the gravitational redshift across a millimetre-scale atomic sample, Nature 602, 420 (2022).
- Akatsuka et al. [2008] T. Akatsuka, M. Takamoto, and H. Katori, Optical lattice clocks with non-interacting bosons and fermions, Nature Physics 4, 954 (2008).
- Campbell et al. [2017] S. L. Campbell, R. B. Hutson, G. E. Marti, A. Goban, N. Darkwah Oppong, R. L. McNally, L. Sonderhouse, J. M. Robinson, W. Zhang, B. J. Bloom, and J. Ye, A fermi-degenerate three-dimensional optical lattice clock, Science 358, 90 (2017).
- Swallows et al. [2011] M. D. Swallows, M. Bishof, Y. Lin, S. Blatt, M. J. Martin, A. M. Rey, and J. Ye, Suppression of collisional shifts in a strongly interacting lattice clock, Science 331, 1043 (2011).
- Anderson and Kasevich [1998] B. P. Anderson and M. A. Kasevich, Macroscopic quantum interference from atomic tunnel arrays, Science 282, 1686 (1998).
- Paredes et al. [2004] B. Paredes, A. Widera, V. Murg, O. Mandel, S. Fölling, I. Cirac, G. V. Shlyapnikov, T. W. Hänsch, and I. Bloch, Tonks-Girardeau gas of ultracold atoms in an optical lattice, Nature 429, 277 (2004).
- Kinoshita et al. [2004] T. Kinoshita, T. Wenger, and D. S. Weiss, Observation of a one-dimensional Tonks-Girardeau gas, Science 305, 1125 (2004).
- Greiner et al. [2002] M. Greiner, O. Mandel, T. Esslinger, T. W. Hänsch, and I. Bloch, Quantum phase transition from a superfluid to a Mott insulator in a gas of ultracold atoms, Nature 415, 39 (2002).
- Bakr et al. [2009] W. S. Bakr, J. I. Gillen, A. Peng, S. Fölling, and M. Greiner, A quantum gas microscope for detecting single atoms in a Hubbard-regime optical lattice, Nature 462, 74 (2009).
- Sherson et al. [2010] J. F. Sherson, C. Weitenberg, M. Endres, M. Cheneau, I. Bloch, and S. Kuhr, Single-atom-resolved fluorescence imaging of an atomic Mott insulator, Nature 467, 68 (2010).
- Cheuk et al. [2015] L. W. Cheuk, M. A. Nichols, M. Okan, T. Gersdorf, V. V. Ramasesh, W. S. Bakr, T. Lompe, and M. W. Zwierlein, Quantum-gas microscope for fermionic atoms, Physical Review Letters 114, 193001 (2015).
- Haller et al. [2015] E. Haller, J. Hudson, A. Kelly, D. A. Cotta, B. Peaudecerf, G. D. Bruce, and S. Kuhr, Single-atom imaging of fermions in a quantum-gas microscope, Nature Physics 11, 738 (2015).
- Parsons et al. [2015] M. F. Parsons, F. Huber, A. Mazurenko, C. S. Chiu, W. Setiawan, K. Wooley-Brown, S. Blatt, and M. Greiner, Site-resolved imaging of fermionic 6Li in an optical lattice, Physical Review Letters 114, 213002 (2015).
- Yamamoto et al. [2016] R. Yamamoto, J. Kobayashi, T. Kuno, K. Kato, and Y. Takahashi, An ytterbium quantum gas microscope with narrow-line laser cooling, New Journal of Physics 18, 023016 (2016).
- Buob et al. [2023] S. Buob, J. Höschele, V. Makhalov, A. Rubio-Abadal, and L. Tarruell, A strontium quantum-gas microscope (2023).
- Su et al. [2023] L. Su, A. Douglas, M. Szurek, R. Groth, S. F. Ozturk, A. Krahn, A. H. Hébert, G. A. Phelps, S. Ebadi, S. Dickerson, F. Ferlaino, O. Marković, and M. Greiner, Dipolar quantum solids emerging in a Hubbard quantum simulator, Nature 622, 724 (2023).
- Sohmen et al. [2023] M. Sohmen, M. J. Mark, M. Greiner, and F. Ferlaino, A ship-in-a-bottle quantum gas microscope setup for magnetic mixtures, SciPost Physics 15, 182 (2023).
- Young et al. [2022] A. W. Young, W. J. Eckner, N. Schine, A. M. Childs, and A. M. Kaufman, Tweezer-programmable 2D quantum walks in a Hubbard-regime lattice, Science 377, 885 (2022).
- Ebadi et al. [2022] S. Ebadi, A. Keesling, M. Cain, T. T. Wang, H. Levine, D. Bluvstein, G. Semeghini, A. Omran, J.-G. Liu, R. Samajdar, X.-Z. Luo, B. Nash, X. Gao, B. Barak, E. Farhi, S. Sachdev, N. Gemelke, L. Zhou, S. Choi, H. Pichler, S.-T. Wang, M. Greiner, V. Vuletić, and M. D. Lukin, Quantum optimization of maximum independent set using Rydberg atom arrays, Science 376, 1209 (2022).
- Graham et al. [2022] T. M. Graham, Y. Song, J. Scott, C. Poole, L. Phuttitarn, K. Jooya, P. Eichler, X. Jiang, A. Marra, B. Grinkemeyer, M. Kwon, M. Ebert, J. Cherek, M. T. Lichtman, M. Gillette, J. Gilbert, D. Bowman, T. Ballance, C. Campbell, E. D. Dahl, O. Crawford, N. S. Blunt, B. Rogers, T. Noel, and M. Saffman, Multi-qubit entanglement and algorithms on a neutral-atom quantum computer, Nature 604, 457 (2022).
- Raab et al. [1987] E. L. Raab, M. Prentiss, A. Cable, S. Chu, and D. E. Pritchard, Trapping of neutral sodium atoms with radiation pressure, Physical Review Letters 59, 2631 (1987).
- Note [1] Often with frequency shifts between non-colinear beams.
- Hopkins and Durrant [1997] S. A. Hopkins and A. V. Durrant, Parameters for polarization gradients in three-dimensional electromagnetic standing waves, Physical Review A 56, 4012 (1997).
- Gallego et al. [2009] D. Gallego, S. Hofferberth, T. Schumm, P. Krüger, and J. Schmiedmayer, Optical lattice on an atom chip, Optics Letters 34, 3463 (2009).
- Straatsma et al. [2015] C. J. E. Straatsma, M. K. Ivory, J. Duggan, J. Ramirez-Serrano, D. Z. Anderson, and E. A. Salim, On-chip optical lattice for cold atom experiments, Optics Letters 40, 3368 (2015).
- Lee et al. [1996] K. I. Lee, J. A. Kim, H. R. Noh, and W. Jhe, Single-beam atom trap in a pyramidal and conical hollow mirror, Optics Letters 21, 1177 (1996).
- Bowden et al. [2019] W. Bowden, R. Hobson, I. R. Hill, A. Vianello, M. Schioppo, A. Silva, H. S. Margolis, P. E. G. Baird, and P. Gill, A pyramid MOT with integrated optical cavities as a cold atom platform for an optical lattice clock, Scientific Reports 9, 11704 (2019).
- Bodart et al. [2010] Q. Bodart, S. Merlet, N. Malossi, F. P. Dos Santos, P. Bouyer, and A. Landragin, A cold atom pyramidal gravimeter with a single laser beam, Applied Physics Letters 96, 134101 (2010).
- Vangeleyn et al. [2009] M. Vangeleyn, P. F. Griffin, E. Riis, and A. S. Arnold, Single-laser, one beam, tetrahedral magneto-optical trap, Optics Express 17, 13601 (2009).
- Nshii et al. [2013] C. C. Nshii, M. Vangeleyn, J. P. Cotter, P. F. Griffin, E. A. Hinds, C. N. Ironside, P. See, A. G. Sinclair, E. Riis, and A. S. Arnold, A surface-patterned chip as a strong source of ultracold atoms for quantum technologies, Nature Nanotechnology 8, 321 (2013).
- McGilligan et al. [2017] J. P. McGilligan, P. F. Griffin, R. Elvin, S. J. Ingleby, E. Riis, and A. S. Arnold, Grating chips for quantum technologies, Scientific Reports 7, 384 (2017).
- Barker et al. [2022] D. S. Barker, E. B. Norrgard, N. N. Klimov, J. A. Fedchak, J. Scherschligt, and S. Eckel, -enhanced gray molasses in a tetrahedral laser beam geometry, Optics Express 30, 9959 (2022).
- Sitaram et al. [2020] A. Sitaram, P. K. Elgee, G. K. Campbell, N. N. Klimov, S. Eckel, and D. S. Barker, Confinement of an alkaline-earth element in a grating magneto-optical trap, Review of Scientific Instruments 91, 103202 (2020).
- Bondza et al. [2022] S. Bondza, C. Lisdat, S. Kroker, and T. Leopold, Two-color grating magneto-optical trap for narrow-line laser cooling, Physical Review Applied 17, 044002 (2022).
- Burrow et al. [2023] O. S. Burrow, R. J. Fasano, W. Brand, M. W. Wright, W. Li, A. D. Ludlow, E. Riis, P. F. Griffin, and A. S. Arnold, Optimal binary gratings for multi-wavelength magneto-optical traps, Optics Express 31, 40871 (2023).
- McGilligan et al. [2020] J. P. McGilligan, K. R. Moore, A. Dellis, G. D. Martinez, E. de Clercq, P. F. Griffin, A. S. Arnold, E. Riis, R. Boudot, and J. Kitching, Laser cooling in a chip-scale platform, Applied Physics Letters 117, 054001 (2020).
- Lee et al. [2022] J. Lee, R. Ding, J. Christensen, R. R. Rosenthal, A. Ison, D. P. Gillund, D. Bossert, K. H. Fuerschbach, W. Kindel, P. S. Finnegan, J. R. Wendt, M. Gehl, A. Kodigala, H. McGuinness, C. A. Walker, S. A. Kemme, A. Lentine, G. Biedermann, and P. D. D. Schwindt, A compact cold-atom interferometer with a high data-rate grating magneto-optical trap and a photonic-integrated-circuit-compatible laser system, Nature Communications 13, 5131 (2022).
- Abend et al. [2023] S. Abend, B. Allard, A. S. Arnold, T. Ban, L. Barry, B. Battelier, A. Bawamia, Q. Beaufils, S. Bernon, A. Bertoldi, A. Bonnin, P. Bouyer, A. Bresson, O. S. Burrow, B. Canuel, B. Desruelle, G. Drougakis, R. Forsberg, N. Gaaloul, A. Gauguet, M. Gersemann, P. F. Griffin, H. Heine, V. A. Henderson, W. Herr, S. Kanthak, M. Krutzik, M. D. Lachmann, R. Lammegger, W. Magnes, G. Mileti, M. W. Mitchell, S. Mottini, D. Papazoglou, F. Pereira dos Santos, A. Peters, E. Rasel, E. Riis, C. Schubert, S. T. Seidel, G. M. Tino, M. Van Den Bossche, W. von Klitzing, A. Wicht, M. Witkowski, N. Zahzam, and M. Zawada, Technology roadmap for cold-atoms based quantum inertial sensor in space, AVS Quantum Science 5, 019201 (2023).
- Cotter et al. [2016] J. P. Cotter, J. P. McGilligan, P. F. Griffin, I. M. Rabey, K. Docherty, E. Riis, A. S. Arnold, and E. A. Hinds, Design and fabrication of diffractive atom chips for laser cooling and trapping, Applied Physics B 122, 172 (2016).
- McGilligan et al. [2016] J. P. McGilligan, P. F. Griffin, E. Riis, and A. S. Arnold, Diffraction-grating characterization for cold-atom experiments, Journal of the Optical Society of America B 33, 1271 (2016).
- McGilligan et al. [2015] J. P. McGilligan, P. F. Griffin, E. Riis, and A. S. Arnold, Phase-space properties of magneto-optical traps utilising micro-fabricated gratings, Optics Express 23, 8948 (2015).
- Webber [2004] M. J. Webber, Handbook of Optical Materials (CRC Press, 2004).
- Squires et al. [2016] M. B. Squires, S. E. Olson, B. Kasch, J. A. Stickney, C. J. Erickson, J. A. R. Crow, E. J. Carlson, and J. H. Burke, Ex vacuo atom chip Bose-Einstein condensate, Applied Physics Letters 109, 264101 (2016).
- Note [2] The lattices here could in principle be achieved with an ex-vacuo grating chip, near a thin-walled glass chamber, but this could reduce the lattice capture efficiency.
- Note [3] Additionally, a notch filter with a peak transmission centred around nm is used when investigating the 1D nm lattice.
- Thompson and Scholten [2012] D. J. Thompson and R. E. Scholten, Narrow linewidth tunable external cavity diode laser using wide bandwidth filter, Review of Scientific Instruments 83, 023107 (2012).
- Daffurn et al. [2021] A. Daffurn, R. F. Offer, and A. S. Arnold, A simple, powerful diode laser system for atomic physics, Applied Optics 60, 5832 (2021).
- Note [4] Alan Bregazzi, PhD thesis (Strathclyde, 2024).
- Arpornthip et al. [2012] T. Arpornthip, C. A. Sackett, and K. J. Hughes, Vacuum-pressure measurement using a magneto-optical trap, Physical Review A 85, 033420 (2012).
- Moore et al. [2015] R. W. G. Moore, L. A. Lee, E. A. Findlay, L. Torralbo-Campo, G. D. Bruce, and D. Cassettari, Measurement of vacuum pressure with a magneto-optical trap: A pressure-rise method, Review of Scientific Instruments 86, 093108 (2015).
- Burrow et al. [2021] O. S. Burrow, P. F. Osborn, E. Boughton, F. Mirando, D. P. Burt, P. F. Griffin, A. S. Arnold, and E. Riis, Stand-alone vacuum cell for compact ultracold quantum technologies, Applied Physics Letters 119, 124002 (2021).
- Wu et al. [2006] J. Wu, R. Newell, M. Hausmann, D. J. Vieira, and X. Zhao, Loading dynamics of optical trap and parametric excitation resonances of trapped atoms, Journal of Applied Physics 100, 054903 (2006).
- Friebel et al. [1998] S. Friebel, C. D’Andrea, J. Walz, M. Weitz, and T. W. Hänsch, CO2-laser optical lattice with cold rubidium atoms, Physical Review A 57, R20 (1998).
- Note [5] Required for balanced diffraction from all three sectors of the grating.
- Note [6] Similar discrepancies between theory and experiment have been reported previously, with experimental frequencies (60-90)% below the maximum theoretical expectations in several separate experiments Ref. [58].
- Struck et al. [2011] J. Struck, C. Ölschläger, R. Le Targat, P. Soltan-Panahi, A. Eckardt, M. Lewenstein, P. Windpassinger, and K. Sengstock, Quantum simulation of frustrated classical magnetism in triangular optical lattices, Science 333, 996 (2011).
- Xu et al. [2023] M. Xu, L. H. Kendrick, A. Kale, Y. Gang, G. Ji, R. T. Scalettar, M. Lebrat, and M. Greiner, Frustration- and doping-induced magnetism in a Fermi-Hubbard simulator, Nature 620, 971 (2023).
- Mongkolkiattichai et al. [2023] J. Mongkolkiattichai, L. Liu, D. Garwood, J. Yang, and P. Schauss, Quantum gas microscopy of fermionic triangular-lattice Mott insulators, Physical Review A 108, L061301 (2023).
- Imhof et al. [2017] E. Imhof, B. K. Stuhl, B. Kasch, B. Kroese, S. E. Olson, and M. B. Squires, Two-dimensional grating magneto-optical trap, Physical Review A 96, 033636 (2017).
- Viebahn et al. [2019] K. Viebahn, M. Sbroscia, E. Carter, J.-C. Yu, and U. Schneider, Matter-wave diffraction from a quasicrystalline optical lattice, Physical Review Letters 122, 110404 (2019).
- Wintersperger et al. [2020] K. Wintersperger, C. Braun, F. N. Ünal, A. Eckardt, M. D. Liberto, N. Goldman, I. Bloch, and M. Aidelsburger, Realization of an anomalous Floquet topological system with ultracold atoms, Nature Physics 16, 1058 (2020).
- Leung et al. [2020] T.-H. Leung, M. N. Schwarz, S.-W. Chang, C. D. Brown, G. Unnikrishnan, and D. Stamper-Kurn, Interaction-enhanced group velocity of bosons in the flat band of an optical kagome lattice, Physical Review Letters 125, 133001 (2020).
- Note [7] TBD.

### v2

@@
 Theoretical cross-sections of the 3D lattice itself are shown in Fig.
-1 a and b, illustrating the different lattice regions with their distinct geometries and dimensions.
-The 1D lattice model is shown in Fig.
-1 c, with experimental data in d.
+1 b and c, illustrating the different lattice regions with their distinct geometries and dimensions.
+The 1D lattice model and experimental data are shown in Fig.
+SM2 a and b, respectively.
 To make the 3D (1D) optical lattices in this paper, we use a single circularly- (linearly-) polarised laser beam, normally incident near the centre of the grating with wavelength nm (nm), power W (W), and intensity radius mm (m).
@@
 The typical MOT location and size are also shown in Fig.
-1 a and b.
+1 b and c.
 The grating chip is bonded to the open end facet of a rectangular cuvette vacuum cell using UHV compatible epoxy, allowing a base vacuum pressure at the mbar level, as measured by the ion pump.
@@
 For both the 3D and 1D lattice geometries s long resonant fluorescent images are taken with the MOT light after a ms time of flight following the extinction of the lattice beam.
-A calculation of atomic motion during imaging (see SM) quantitatively matches the large measured width of the cloud seen in Fig.
-1 d.
 In both the 3D and 1D lattices, the lattice beam power can be rapidly adjusted by a single-pass acousto-optic modulator (AOM) designed for the beam wavelength.
@@
 A log plot of the number of trapped atoms as a function of hold time in both the 3D and 1D lattice is shown in Fig.
-1 e.
+2.
 Exponential decays are fitted, excluding data where ms, to eliminate untrapped atoms ejected from the imaging region over this time scale due to gravity.
@@
 By plotting the primary observed resonance frequency as a function of , resonance data for different 3D lattice trap depths can be summarised in a single plot, as shown in Fig.
-2b.
+3b.
 As expected, these data follow a relation, where here the fitted value is kHzGHz/mW, in good agreement with theory.
@@
 This is due to lower spatial overlap of the 1D lattice with the MOT and hence lower initial atom number (Fig.
-1 e), as well as rapid radial expansion during fluorescence imaging (Fig.
-1 c, d and the SM).
+2), as well as rapid radial expansion during fluorescence imaging (Fig.
+SM2 and discussion in the SM).
 The 1D lattice beam focal position on the grating surface was optimised for atom number, and given the lattice beam Rayleigh range of mm, the beam waist was constant over the entire lattice.
@@
 However, we note that by loading the MOT atoms into a magnetic and/or optical trap, they can be transported beyond the MOT capture volume, and thus future single-input-beam two-dimensional lattices are within easy reach (Fig.
-1 a, b).
+1 b, c).
 We have demonstrated experimentally that the GMOT architecture can be extended from a single-input-beam MOT, to both three- and one-dimensional single-input-beam optical lattices on the same chip that is also likely to enable 2D lattices.
@@
 For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising from this submission.
-The dataset is available here 777TBD.References
+The dataset is available here 777Dataset DOI to be added., and supplementary material which includes Refs.
+[70, 71, 72] can be found here 888Supplementary Material link to be added.References
 - Grimm et al. [2000] R. Grimm, M. Weidemüller, and Y. B. Ovchinnikov, Optical dipole traps for neutral atoms, in Advances In Atomic, Molecular, and Optical Physics (Elsevier, 2000) pp. 95–170.
@@
 Stamper-Kurn, Interaction-enhanced group velocity of bosons in the flat band of an optical kagome lattice, Physical Review Letters 125, 133001 (2020).
-- Note [7] TBD.
+- Note [7] Dataset DOI to be added.
+- Vangeleyn et al. [2010] M. Vangeleyn, P. F. Griffin, E. Riis, and A. S. Arnold, Laser cooling with a single laser beam and a planar diffractor, Optics Letters 35, 3453 (2010).
+- Henderson et al.
+[2020] V.
+A.
+Henderson, M.
+Y.
+H.
+Johnson, Y.
+B.
+Kale, P.
+F.
+Griffin, E.
+Riis, and A.
+S.
+Arnold, Optical characterisation of micro-fabricated Fresnel zone plates for atomic waveguides, Optics Express 28, 9072 (2020).
+- Arnold and Manson [2000] A. S. Arnold and P. J. Manson, Atomic density and temperature distributions in magneto-optical traps, Journal of the Optical Society of America B 17, 497 (2000).
+- Note [8] Supplementary Material link to be added.